# Latent learning: episodic memory complements parametric learning by enabling flexible reuse of experiences

### v2

@@ 3 Towards a formalization of latent learning
We would like to distinguish this from many types of generalization that a sufficiently expressive and implicitly/explicitly regularized model will acquire from parametric learning over sufficient data. First, the model will learn transferable knowledge about the tasks: if there is a systematic input-output relationship for task $t$, models will tend to learn the mappings in a way that generalizes $t$ to novel inputs that provide sufficient information in context (subject to the usual constraints on generalization in machine learning, e.g., under distribution shift). Second, the model will learn simple associations among the inputs, task cues, and outputs. In some cases, this may allow improved performance even on input, task pairings $x^{\prime},t^{\prime}$ not seen in training; as a trivial example if for a given input $x$ the output is always constant ($\forall t\in T\;f(x,t)=c$), the model will not need to see all possible tasks to learn this. Similarly if multiple task cues are redundant, if certain tasks only have a subset of possible outputs, etc., the model may readily learn these regularities.
@@
  • Reversal without context $\bm{[x^{\prime},t^{\prime},f(x,t^{\prime})]}$: E.g., “Who taught Aristotle?” “Plato.” The crucial test of latent learning. The necessary information has been learned, and the question could be answered with that information in context (the two conditions above). However, we claim that the model will generally not answer this condition correctly from parametric learning alone.
@@
Note that this description is intended to provide formal intuitions, not to fully characterize all the forms in which a latent learning challenge may occur in practice; for example, in language modeling there is usually no complete disentanglement between the observations $x$, the task indicators $t$ and the outputs $f$ to be learned. This entanglement also occurs in some of our examples, e.g., it is difficult to fully separate what is observations from task indicators in the gridworld environment. Likewise, there are other forms of redundancy in $f$ (aside from non-injectivity), that could enable latent generalization. We leave fully characterizing the extent of latent learning to future work.
@@ 4.3 Semantic structure dataset
We also create two versions of the dataset; one in which there are strong associative cues that can provide shortcuts to inferring an answer (e.g., if all birds have wings, even if the statement “eagles have wings” is held out, the associations between “eagle” and “bird” may be sufficient to infer it), and another in which the associative cues are reduced, by choosing the tests and the distractor choices from the same branch of the tree (so that they share similar associative features). We use these two versions to show how associative cues affect generalization.
@@ 6 Results
(a) With strong associative cues.
(b) Reduced associative cues.
Figure 5: Results on the semantic structure benchmark, comparing performance of a baseline model and one with oracle retrieval, in settings with and without strong associative cues. (5(a)) When strong associative cues are present in the data, both models achieve relatively high performance due to the possibility of associative learning. This demonstrates how associations can provide an alternative route to generalization in some cases. (5(b)) When associative cues are reduced, the advantage of the retrieval model is more notable. However, in all cases the benefits of retrieval are more muted than in other cases, likely because there are not sufficient examples for the model to acquire strong ICL capabilities.
@@
Adding retrieval:  In Fig. 4, we show that oracle retrieval could solve the latent learning problems described above. On both the Simple Reversals and Codebooks benchmarks, a system equipped with retrieval not only solves the same validation tasks as the baseline model, but also generalizes very well to the key latent test conditions. This shows the striking difference in parametric and non-parametric generalization on the latent learning tests.
Semantic structure benchmark:  We next explore two versions of the more complex semantic structure benchmark (Fig. 5). These experiments demonstrate several interesting effects. First, in the presence of strong associative cues, even the baseline model can generalize fairly well to holdouts like reversals. These associative cues are due to, for example, the fact that other entities similar to the one in question tend to have similar attributes. When these associative cues are removed from the dataset, however, performance of both models drops—and the benefits of retrieval become clearer. However, the benefits of retrieval still remain limited, presumably because this dataset lacks within-document ICL examples from which the model could learn how to most effectively use memory (see below).
@@
Within-example ICL helps learn to use retrieved memories:  We now highlight one result on how ICL interacts with retrieval. Specifically, we find that having training examples that promote within-document in-context learning (ICL) is important for achieving strong benefits of episodic retrieval (Fig. 7). To do so, we create a variation of the simple reversals dataset that omits the ICL-like sequences (i.e., the ones where a forward and backward sequence are included within the same document). Although the oracle retrieval model still experiences retrieved sequences that could be used to learn how to use memory, it in fact fails to achieve strong performance on the latent test split (though it does achieve non-zero performance, unlike the baseline model). These results show that learning to use context within a training example can support learning to use effective retrieval across examples. This finding may explain the small benefits of retrieval on the Semantic Structure benchmark—where there are not clean ICL examples.
@@ 7 Discussion
On generalization from parametric learning:  As noted above, we are absolutely not arguing that models do not generalize out of distribution from parametric learning. There is substantial evidence that models can and will generalize—even out of distribution—under controlled tests (e.g., abreu2025taxonomy). Our results on the Semantic Structure benchmark illustrate one (associative) mechanism by which certain kinds of generalization to systematic holdouts is possible. Indeed, this kind of generalization is one of the core phenomena that motivated the original neural network models of semantic cognition (rogers2004semantic). However, there has also been ample evidence of stronger structural generalization from deep learning models—even in areas that have been considered challenges, such as compositional or syntactic generalization (e.g., murty2023grokking; misra2024language), or generalizing to novel causal structures (lampinen2023passive). Thus, our claim here is simply that we are identifying one particular mode of generalization that appears to be inadequately achieved by gradient-based parametric learning alone in present architectures, and which can be more effectively supported through episodic memory and retrieval.
@@
However, as noted above, other works have found instances of surprising generalization from language models in similar settings to ours (e.g., berglund2023taken; cook2025programming; meinke2023tell). For example, berglund2023taken show that models can generalize from training facts (in separate documents) like “The Pangolin chatbot answers in German” and “Latent AI makes Pangolin” to answer in German when prompted to answer as “Latent’s AI.” This finding may seem to show a type of latent learning that contradicts our claims above. However, in this case there may also be associative cues that are sufficient to give away the answer—even if it requires multiple steps of association (note that we observe some associative multi-hop generalization on the syllogisms in the semantic structure benchmark). Moreover, the absolute degree of generalization observed in these studies is usually relatively low. Nevertheless, we believe that fully exploring the boundary between the types of generalization explored in this paper and other prior works is an important direction for future work. Factors like data diversity (along both relevant and orthogonal axes of variation) likely influence the exact structure of generalization.
@@
Complementary learning systems:  In motivating this work, we highlighted that complementary learning systems theory (mcclelland1995there; kumaran2016learning) emphasizes that episodic retrieval and parametric knowledge play complementary roles in natural intelligence. While past work in this area has often emphasized the benefits of episodic memory for avoiding catastrophic interference or forgetting, our results highlight a different perspective; that episodic retrieval may unlock a fundamentally more flexible route to using information from past learning experiences, even if that information was not relevant to the task at learning time.
@@
Multiplicity of solutions & online vs. offline replay:  We want to highlight that test-time (online) episodic retrieval is not the only solution to the types of problems that we explore. For example, several works have explored training-time (offline) augmentation to address challenges like reversals or multi-hop inferences (akyurek2024deductive; lampinen2025generalization; yang2025synthetic). Other work on “preplaying” routes towards possible alternative goals while learning to navigate a new environment (carvalho2025preemptive) similarly highlights ways to enrich learning such that future generalization will be improved. These training-time solutions require more inference at training time, but via this amortization avoid the retrieval problem at test time. On the other hand, these approaches also require estimating ahead of time which alternative goals may be relevant; retrieval is more flexible. Thus, we expect that training-time augmentation methods likely complement test-time retrieval ones; ultimately natural intelligence is likely to rely on many strategies to improve its chances of generalization.
Indeed, there is substantial evidence that in natural intelligence, hippocampal replay occurs both online and offline, and plays multiple roles (comrie2024hippocampal; eldar2020roles). Online, replay helps to make predictions (kay2020constant) and support flexible decision making (eldar2020roles); offline, replay contributes to consolidation and generalization (kumaran2012generalization; momennejad2018offline; liu2019human). We suspect that artificial intelligence will similarly benefit from employing multiple complementary approaches. Our goal here is simply to highlight that episodic memory retrieval can play an important role in various types of generalization, by bringing learning experiences back into context, where they can be used more flexibly to make inferences for the present or the future.
@@ Acknowledgements
We thank Sridhar Thiagarajan and Jörg Bornschein for helpful comments and suggestions. We thank Murray Shanahan for support.

### v3

@@ 3 Towards a formal framework for latent learning
We would like to distinguish this from many types of generalization that a sufficiently expressive and implicitly/explicitly regularized model will acquire from parametric learning over sufficient data. First, the model will learn transferable procedural knowledge about the tasks: if there is a systematic input-output relationship for task $t$, models will tend to learn the mappings in a way that generalizes $t$ to novel inputs that provide sufficient information in context (subject to the usual constraints on generalization in machine learning, e.g., under distribution shift). Second, the model will learn the associations and similarity relationships among the inputs, task cues, and outputs. In some cases, this may allow improved performance even on input, task pairings $x^{\prime},t^{\prime}$ not seen in training; as a trivial example if for a given input $x$ the output is always constant ($\forall t\in T\;f(x,t)=c$), the model will not need to see all possible tasks to learn this. Similarly if multiple task cues are redundant, if certain tasks only have a subset of possible outputs, etc., the model may readily learn these regularities.
@@
  • Reversal without context $\bm{[x^{\prime},t^{\prime},f(x,t^{\prime})]}$: E.g., “Who taught Aristotle?” “Plato.” The crucial test of latent learning. The necessary information has been learned, and the question could be answered with that information in context (the two conditions above). However, we claim that a causal [^2] Of course, the reversal failure is specific to causal language models; models that do bidirectional prediction over the same data would not exhibit it (wu2024exploring). This finding is compatible with the perspective expressed here, because changing the prediction task to be bidirectional makes the reversal information more explicit in the training task. This observation illustrates the broader point that what is “explicit” or “latent” in a dataset is not purely a function of the data, but also the architecture and objective used to train it. Nevertheless, we hypothesize that bidirectional models would still exhibit the other, more complex latent learning challenges we describe, e.g. in the next paragraph. language model will generally not answer this condition correctly from parametric learning alone.
@@
As noted above, this formal is more limited than the full scope of settings in which latent learning challenges occur; for example, in language modeling there is usually no complete disentanglement between the observations $x$, the task indicators $t$ and the outputs $f$ to be learned. This entanglement also occurs in some of our examples, e.g., it is difficult to fully separate observations from task indicators in the gridworld environment. Likewise, there are other forms of redundancy in $f$ (aside from non-injectivity), that could enable latent generalization. We leave fully characterizing the extent of latent learning to future work. However, we hope that this formal statement makes it more clear where systems might suffer from a challenge.
@@ 4.3 Semantic structure dataset
We also create two versions of the dataset. In the first, there are strong similarity-based cues that can provide shortcuts to inferring an answer. For example, if there many birds—all or most of which share properties like flying, nesting, having wings, having hollow bones, etc.—even if the particular statement “eagles have wings” is held out, the model may be able to infer it via similarity-based generalization from other birds (e.g., rogers2004semantic). We create another version of the dataset in which these similarity-based cues are reduced, by making the semantic structure tree wider and shallower (more heavily branching at the top, so that fewer leaf nodes have common properties) and choosing the test and the distractor choices from the same branch of the tree (so that all the possible choices have many similar features). We use these two versions to show how similarity cues affect parametric and episodic generalization.
@@ 6 Results
(a) With strong similarity-based cues.
(b) Reduced similarity-based cues.
Figure 5: Results on the semantic structure benchmark, comparing performance of a baseline model and one with oracle retrieval, in settings with and without strong associative cues. (5(a)) When strong similarity-based cues are present in the data, both models achieve relatively high performance due to the possibility of associative learning. This demonstrates how similarity-based generalization can provide an alternative route to generalization in some cases. (5(b)) When similarity-based cues are reduced, the advantage of the retrieval model is more notable. However, in all cases the benefits of retrieval are more muted than in other cases, likely because there are not sufficient examples for the model to acquire strong ICL capabilities.
@@
Adding retrieval:  In Fig. 4, we show that oracle retrieval could solve the latent learning problems described above. On both the Simple Reversals and Codebooks benchmarks, a system equipped with retrieval not only solves the same validation tasks as the baseline model, but also generalizes very well to the key latent test conditions. This shows the striking difference in parametric and retrieval-based generalization on the latent learning tests.
Semantic structure benchmark:  We next explore two versions of the more complex semantic structure benchmark (Fig. 5). These experiments demonstrate several interesting effects. First, in the presence of strong similairty-based cues, even the baseline model can generalize fairly well to holdouts like reversals. These similarity cues are due to, for example, the fact that other entities similar to the one in question tend to have similar attributes. When these associative cues are removed from the dataset, however, performance of both models drops—and the benefits of retrieval become clearer. However, the benefits of retrieval still remain limited, presumably because this dataset lacks within-document ICL examples from which the model could learn how to most effectively use memory (see below).
@@
Within-example ICL helps learn to use retrieved memories:  We now highlight one result on how ICL interacts with retrieval. Specifically, we find that having training examples that promote within-document in-context learning (ICL) is important for achieving strong benefits of episodic retrieval (Fig. 7). To do so, we create a variation of the simple reversals dataset that omits the ICL-like sequences (i.e., the ones where a forward and backward sequence are included within the same document). Although the oracle retrieval model still experiences retrieved sequences that could be used to learn how to use memory, it in fact fails to achieve strong performance on the latent test split (though it does achieve non-zero performance, unlike the baseline model). These results show that learning to use context within a training example can support learning to use effective retrieval across examples. This finding aligns with the weak benefits of retrieval on the Semantic Structure benchmark—where there are not clean ICL examples.
@@ 7 Discussion
On generalization from parametric learning:  As noted above, we are absolutely not arguing that models do not generalize out of distribution from parametric learning. Indeed, the flexibility of in-context learning to apply a procedure such as reversal or encoding to an instance that has never been seen in training is driven precisely by the procedural generalization of parametric learning. Our results on the semantic structure benchmark illustrate another mechanism—similarity-based generalization—through which certain kinds of generalization to systematic holdouts is possible, even in cases where latent learning might otherwise be necessary. Modes of generalization like these presumably explain many of the instances of successful generalization of modern machine learning. Thus, our claim here is simply that the type of latent learning we highlight is one particular mode of generalization that appears to be inadequately achieved by gradient-based parametric learning alone in present architectures, and which can be more effectively supported through episodic memory and retrieval.
@@
However, as noted above, other works have found instances of surprising generalization from language models in similar settings to ours (e.g., berglund2023taken; cook2025programming; meinke2023tell). For example, berglund2023taken show that models can generalize from training facts (in separate documents) like “The Pangolin chatbot answers in German” and “Latent AI makes Pangolin” to answer in German when prompted to answer as “Latent’s AI.” This finding may seem to show a type of latent learning that contradicts our claims above. However, in this case there may also be similarity-based or associative cues that are sufficient to give away the answer—even if it requires multiple steps of association (note that we observe some multi-hop similarity-based generalization on the syllogisms in the semantic structure benchmark). Moreover, the absolute degree of generalization observed in these studies is usually relatively low. Nevertheless, we believe that fully exploring the boundary between the types of generalization explored in this paper and other prior works is an important direction for future work. Factors like data diversity (along both relevant and orthogonal axes of variation) likely influence the exact structure of generalization.
@@
Complementary learning systems:  In motivating this work, we highlighted that complementary learning systems theory (mcclelland1995there; kumaran2016learning) emphasizes that episodic retrieval and parametric knowledge play complementary roles in natural intelligence. While past work in this area has often emphasized the benefits of episodic memory for learning from few exposures without interference, and assisting in integrating new information, our results highlight a different perspective; that episodic retrieval may unlock a fundamentally more flexible route to using information from past learning experiences, even if that information was not relevant to the task at learning time.
@@
Multiplicity of solutions & online vs. offline replay:  We want to highlight that test-time (online) episodic retrieval is not the only solution to the types of problems that we explore. For example, several works have explored training-time (offline) augmentation to address challenges like reversals or multi-hop inferences (akyurek2024deductive; lampinen2025generalization; yang2025synthetic). Other work on “preplaying” routes towards possible alternative goals while learning to navigate a new environment (carvalho2025preemptive) similarly highlights ways to enrich learning such that future generalization will be improved. These solutions can be interpreted as applying effectively the same computational trick—using episodic retrieval to make latent information available for more flexible reuse—but doing so ahead of time to cache the solution for later. These training-time solutions therefore require more computational effort at learning time, but via this amortization avoid the retrieval problem at test time. On the other hand, these approaches also require estimating ahead of time which alternative goals may be relevant for future use; retrieval is more flexible. Thus, we expect that training-time augmentation methods likely complement test-time retrieval ones; ultimately natural intelligence is likely to rely on many strategies to improve its chances of generalization.
Indeed, there is substantial evidence that in natural intelligence, hippocampal replay occurs both online and offline, and plays multiple roles (comrie2024hippocampal; eldar2020roles; cowan2021memory). Online, replay helps to make predictions (kay2020constant) and support flexible decision making (eldar2020roles); offline, (p)replay contributes to consolidation and generalization (olafsdottir2015hippocampal; kumaran2012generalization; momennejad2018offline; liu2019human). We suspect that artificial intelligence will similarly benefit from employing multiple complementary approaches. Our goal here is simply to highlight that episodic memory retrieval can play an important role in various types of generalization, by bringing learning experiences back into context, where they can be used more flexibly to make inferences for the present or the future.
@@ Acknowledgements
We thank Sridhar Thiagarajan, Jörg Bornschein, and Tyler Bonnen for helpful comments and suggestions. We thank Murray Shanahan for support.